Jeremy has a skateboard ramp sent 30° from the ground if the height of the platform is 8 feet from the ground how long is the ramp up to the top of the platform?
Answer by addingup(3677)

So 30 is the angle formed by the ground and the beginning of the ramp. 8 is the height of the ramp as measured from the ground to the top of the ramp. Thus, we have a right triangle where we know the angle and the opposite side.
To solve it, we use SOH (Sine(theta) Opposite/Hypothenuse)
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Sin(30)= 8/h (use your calculator to find sin(30)
1/2 = 8/h We take the reciprocal of both sides:
2/1 = h/8; 2= h/8 multiply both sides by 8
2*8= h
16 = h which is the length of your ramp.